OpenStudy (anonymous):

tan^-1(sqr3/3) i dont have the calculator for this one.

OpenStudy (anonymous):

30 degrees or pi/6

OpenStudy (anonymous):

i have to find the angles that fit the expression

OpenStudy (anonymous):

Actually, if you have Google, you have a calculator. For example, google arctan (sqrt 3/2) and get arctan(sqrt(3) / 2) = 0.713724379 rad

OpenStudy (anonymous):

And if that doesn't do it, there's always Wolfram Alpha: http://www.wolframalpha.com/

OpenStudy (anonymous):

thank you but i stuck finding the angles that fit the expression

OpenStudy (anonymous):

what expression?

OpenStudy (anonymous):

For that matter, if you're on Windows, you have a built-in scientific calculator that would be more than enough for this. Just choose scientific from the view menu on the standard windows calc.

OpenStudy (anonymous):

i have to find the angle values that fit the expression.

OpenStudy (anonymous):

If you're looking for the angles, then you just run the inverse function to solve for the angle. 

OpenStudy (anonymous):

On Windows, the inverses are sinh, cosh, tanh.

OpenStudy (anonymous):

i got 33.869

OpenStudy (anonymous):

Did you mean to type \[\theta = \tan^{-1} \frac{\sqrt{3}}{3}\]

OpenStudy (anonymous):

Because that give me an even 30 degrees.

OpenStudy (anonymous):

yes that what i ment i dont know how to imput it on here

OpenStudy (anonymous):

Ah, use the equation button. The syntax is called LaTeX, but you don't need to know it to use the editor. So solve in this order: \[\sqrt{3}\] Divide that by 3. Press the arctangent button (tanh, atan or cot, depending on your calculator)

OpenStudy (anonymous):

You should then get the same answer.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem.